### v2

# Nondeterminism-Aware Optimistic Verification for Floating-Point Neural Networks

Abstract.
Neural networks increasingly run on hardware outside the user’s control (cloud GPUs, inference marketplaces, edge specialized accelerators) for both training and inference. Yet ML-as-a-Service reveals little about what actually ran or whether returned outputs faithfully reflect the intended inputs. Users lack recourse against service downgrades such as model swaps, quantization, graph rewrites, or discrepancies like altered advertisement embeddings. Verifying outputs is especially difficult because floating-point execution on heterogeneous accelerators is inherently nondeterministic. Existing approaches like zkML, deterministic replay, TEEs, and replication are either impractical for real floating-point neural networks or reintroduce vendor trust. We present NAO: a Nondeterministic tolerance Aware Optimistic verification protocol for floating-point neural networks that accepts outputs within principled operator-level acceptance regions rather than requiring bitwise equality. NAO combines two complementary error models: (i) sound per-operator IEEE-754 worst-case bounds and (ii) tight empirical percentile profiles calibrated across hardware. Discrepancies are resolved via a Merkle-anchored, threshold-guided interactive dispute game that recursively partitions the traced computation graph until one operator remains; at the leaf, adjudication reduces to either a lightweight theoretical-bound check or a small honest-majority vote against empirical thresholds. Unchallenged results finalize after a challenge window, without requiring trusted hardware or deterministic kernels.
We implement NAO as a PyTorch-compatible runtime and a contract layer currently deployed on Ethereum Holesky testnet. The runtime instruments graphs, computes per-operator bounds in real time, and runs unmodified vendor kernels in FP32 with negligible overhead (0.3% additional latency on Qwen3-8B). Our evaluation spans CNNs, Transformers, LLMs, and diffusion models across multiple GPUs (A100, H100, RTX6000, RTX4090). Results show that empirical thresholds are $10^{2}$$10^{3}\times$ tighter than theoretical bounds, and bound-aware adversarial attacks achieve 0% success under empirical checks. Together, NAO reconciles scalability with verifiability for real-world heterogeneous ML compute.
copyright: none
1. Introduction
Modern machine learning (ML) systems increasingly outsource training and inference to third-party infrastructure, including public clouds, inference marketplaces, edge clusters, and specialized accelerators (Zhang et al., 2019). This outsourcing is driven by the rapidly growing cost of ML computation, which has led to two notable trends. First, to conserve scarce compute budgets, practitioners often trade accuracy for cost at serving time, for example, by switching to smaller models, lowering precision, or coarsening quality knobs. Second, platform opacity makes these trade-offs invisible: in ML-as-a-service (MLaaS) platforms, whether closed-source proprietary models or open-model inference services, users cannot see which model was executed, what inputs were processed, or whether results faithfully reflect the intended computation. As a result, users accept outputs they cannot independently verify, leaving them exposed to silent downgrades such as model swaps, early exit (Lahiany and Aperstein, 2022), or quantization (Zhao et al., 2024), or service discrepancies like altered advertisement embeddings. When consumers of results do not control the underlying hardware, verifiable ML becomes a crucial system problem: the user should be provided with an efficiently verifiable proof certifying that the output was generated by executing an agreed model on agreed inputs (Jia et al., 2021; Choi et al., 2023).
As demand for ML compute grows, platforms with decentralized inference markets and heterogeneous GPUs fleets promise greater capacity and price diversity. However, they also expose a fundamental technical obstacle for verifiable ML: modern accelerators do not guarantee deterministic outputs. Floating-point arithmetic (IEEE-754) (Kahan, 1996) is non-associative; vendor kernels legitimately reorder reductions and fuse operators; thread scheduling and atomics introduce nondeterministic execution traces (Chou et al., 2020; Shanmugavelu et al., 2024). Consequently, running the same neural network (NN) on different hardware, or even repeated executions on the same GPU can yield slightly different outputs. Here, we use operator to denote a primitive tensor-level function (a node in the NN graph), such as matrix multiplication, convolution, element-wise addition, or activation in PyTorch/ATen. In short, cross-platform nondeterminism is intrinsic to production GPU stacks, yet most existing verification methods assume exact reproducibility. This work addresses the gap.
Figure 1. Tolerance-aware optimistic verification on two example tasks. Left: Using LLM agents for automated transaction Honest proposer ($y_{0}$) returns the correct destination: 0xA173****d422, amount: 0.01. Malicious ($y_{1}$) applies unbounded node-wise perturbations and outputs an absurd amount to send (rejected). ($y_{2}$) stays within IEEE-754 worst-case bounds but violates empirical thresholds (slashed). ($y_{3}$) fits both bounds (accepted). Right: Image classification the honest proposer outputs “Cat”, while three similar types of malicious proposers output “Rock” (rejected), “Rabbit” (rejected), and “Cat” (accepted). Middle: tolerance-aware check. Malicious proposers inject perturbations after each operator. Results are rejected if they fall outside IEEE-754 theoretical bounds or empirical thresholds, while bounded, no explicit deviations are accepted. ($\oplus$: post-operator add)
Why existing approaches fall short. (1) zkML: A natural path is to prove every arithmetic operation in zero knowledge. However, current end-to-end zkML pipelines remain orders of magnitude slower and memory-intensive for realistic floating point models (Wang and Hoang, 2023; Chen et al., 2024; Qu et al., 2025; Sun et al., 2024). Expressing large PyTorch graphs in field arithmetic or fixed-point circuits, and proving native FP32 GPU kernels at scale, is beyond today’s practicality. Moreover, real GPU stacks are inherently nondeterministic: one must either fix a specific execution schedule (sacrificing performance and hardware heterogeneity) or encode acceptance intervals, neither of which current zkML frameworks support. (2) Deterministic replay: Another approach is to enforce deterministic libraries, kernels, and reduction orders so that independent re-executions are bitwise reproducible (Zheng et al., 2021). This discards mature vendor libraries, requires extensive manual redesign of parallel reductions, and constrains accelerators (GPUs/TPUs) that are fundamentally nondeterministic by design. (3) Trusted execution environments (TEEs) attest that a computation ran inside an trusted enclave (Tramer and Boneh, [n. d.]; Moon et al., 2025; Bai et al., 2025). But this shifts trust to single-vendor microcode, introduces side-channel risks (Wei et al., 2018) and performance degradation (Niu et al., 2022) due to limited storage and compute capacity. (4) Replication and majority voting: Protocol-level schemes replicate the computation across multiple providers and accept the majority output. The high cost of replication for large models severely limits practicality and scalability (Zhao et al., 2005; Esmaeili and Mtibaa, 2024). In summary, none of these approaches simultaneously accommodate floating-point nondeterminism, preserve production-level performance, and enable portable, vendor-agnostic verification.
Our approach: verify up to principled error thresholds. NN models are inherently robust to small rounding differences, and user-visible semantics rarely depend on bit-exact equality. Thus, the appropriate notion of correctness is approximate: an output is acceptable if it falls within a model- and operator-specific tolerance induced by IEEE-754 arithmetic. The key challenge is to formalize this tolerance precisely, make it compositional across operators, and ensure that it can be verified efficiently.
This paper: NAO. We present NAO, an tolerance-aware optimistic execution protocol for verifiable neural network computation on heterogeneous accelerators. Instead of requiring bitwise equality, NAO verifies results up to model-specific, operator-level acceptance regions-sets of outputs guaranteed to remain within a prescribed error bound. Discrepancies are resolved through a Merkle-anchored interactive dispute game that localizes disagreements to a single operator (a leaf operator in the Merkle tree (Merkle, 1987)). At that leaf, lightweight adjudication suffices, either through a certified theoretical bound check or a small committee vote against pre-calibrated empirical thresholds. The key idea is to embrace, rather than eliminate, floating-point nondeterminism by verifying neural outputs up to principled per-operator error bounds, while preserving native GPU performance and hardware heterogeneity. By pushing verification down to the granularity where it is both meaningful and efficient: NAO reconciles scalability with accountability. Users can run the models they want, on any hardware the market provides, without compromising correctness, transparency, or recourse.
Two complementary error models. Even when software determinism is enforced (e.g., fixed kernel choices, disabled autotuning), parallel execution still reorders operations, introducing nondeterminism due to the non-associativity of IEEE-754 arithmetic. Rather than attempting to eliminate this through brittle execution constraints, we model and bound rounding error directly. NAO introduces two complementary error models that make tolerance-aware verification both principled and deployable:
$\bullet$ Theoretical IEEE-754 bounds. For each operator, we compute a worst-case element-wise rounding error as a function of its inputs. This bound is sound by construction but overly conservative in deep networks, so we apply it only at the leaf level, where it remains both cheap and decisive.
$\bullet$ Empirical error percentile thresholds. We calibrate rounding errors offline across accelerators to capture the distribution of cross-hardware deviations for each operator. These empirical profiles are tight, model-specific, scalable, and only needed to be calibrated once (easy to post-verify). We use them to guide the search for the disputed leaf and, when necessary, to adjudicate with a small honest-majority committee, where they are more expensive than theoretical bounds but also significantly tighter and more robust.
Our traced runtime instruments PyTorch graphs (Ansel et al., 2024), computes theoretical bounds on the fly, and records intermediate results to build robust cross-device empirical percentile profiles. Instead of propagating worst-case error across the entire graph, we turn composition into localization: disagreements are recursively narrowed until a single operator can be certified. By combining both error bounds, NAO flexibly balances universality, tightness, and scalability.
From whole-model proofs to single-operator checks. Verifying a large model end-to-end against any bound remains expensive. NAO makes verification cheap by localizing disagreement and reducing the verification burden. Model owners provide cryptographic commitments to the weights and graph topology. In the common case, a proposer (compute provider) runs the model and posts a commitment; if unchallenged within a window, the commitment attains finality. If challenged, an interactive $N$-way partition protocol is invoked. At each round, empirical thresholds guide the split: the proposer partitions the traced graph into disjoint subgraphs with inclusion proofs, and the challenger identifies the first child whose outputs exceed its empirical thresholds. After $O(\log_{N}|V|)$ rounds ($V$ is the set of operators), the interactive dispute game narrows to a single operator with agreed inputs, where a theoretical-bound check or a committee vote evaluation resolves the dispute.
Coordination layer. The protocol needs a coordination service that (a) stores tamper-evident commitments to model, graph, and outputs; (b) manages challenge windows and state transitions; and (c) escrows bonds and enforces rewards and punishments. One straightforward option is to deploy a trusted centralized coordinator. In our prototype, however, we instantiate this layer on a decentralized ledger (Ethereum), which provides authenticated logs, fair timeouts, and bond management out of the box, while also offering a transparent, widely understood cost model. Importantly, NAO itself does not rely on blockchain-specific assumptions.
Open-model setting and service transparency. NAO targets the prevalent and rapidly growing open-model setting, where both weights and graph structure are public (e.g., community LLMs). This transparency enables subgraph extraction, per-operator thresholding, and provenance commitments. Closed-model APIs can also benefit, if providers expose such commitments to permissioned parties.
Adversary and attack coverage. Tolerance-aware verification raises two concerns: trustworthiness and attack surface given the error bounds. We adopt a strong threat model in which the proposer may be malicious, injecting perturbations into intermediate tensors to evade empirical thresholds or theoretical error bound checks. To evaluate this risk, we design adaptive gradient-based attacks (Madry et al., 2018): distributional evasion against empirical envelopes and element-wise evasion against theoretical bounds. We further study whether many small, locally admissible deviations could accumulate into a task-level error without being detected. Our results show that the admissible per-operator perturbations are small and do not naturally compound, making end-to-end attacks under NAO practically difficult. Example is in Fig. 1.
Implementation overview. We implement NAO as a PyTorch runtime and a set of Ethereum smart contracts. The PyTorch runtime instruments computation graphs, computes per-operator bounds on the fly, records intermediate traces, partitions and re-executes subgraphs, and emits or verifies Merkle commitments. The smart contracts manage commitments, dispute state transitions, per-round bonds, and slashing. The runtime leverages software-level determinism features, runs unmodified vendor kernels in FP32, and uses FP64 arithmetic for error-bound calculations. It introduces negligible overhead in the optimistic execution ($\approx$0.3% additional latency on Qwen3-8B) and requires no extra memory beyond native subgraph execution (Sec. 7).
Evaluation highlights. Across four representative models (ResNet-152, BERT-large, Qwen3-8B, Stable Diffusion v1-5) and four GPUs (RTX 4090, RTX 6000, A100, H100), we find empirical per-operator errors are $10^{2}\text{--}10^{3}\times$ tighter than worst-case IEEE 754 theoretical bounds for transformers (Fig. 6). Under meticulously designed attacks, empirical error percentile thresholds yield a $0\%$ attack success rate (ASR) and $0\%$ false positives across models, even when thresholds are relaxed by a factor of $3$. By contrast, relying solely on theoretical bounds leaves a small but non-negligible gap: up to $2.4\%$ ASR on Qwen3-8B(Table 2). These results confirm that theoretical worst-case bounds can be overly loose, while empirical thresholds are both tight and robust, motivating committee-based adjudication at the leaf when needed.
Our contributions.
  1. (1)
    Semantics for economically verifiable floating-point ML. We formalize tolerance-aware correctness for tensor programs and design an optimistic challenge-response protocol that achieves economic finality without requiring determinism.
  2. (2)
    Error analysis and adversarial attacks at operator granularity. We develop dual error models: portable theoretical IEEE-754 bounds and tight empirical percentile profiles, and conduct a comprehensive study of bound-aware adversarial attacks, quantifying their limited ability to subvert NAO across representative models.
  3. (3)
    Efficient and robust dispute resolution. We design a Merkle-anchored, threshold-guided dispute game that recursively localizes disagreements to a single operator, supports verifiable subgraph provenance, and enables low-cost adjudication that respects nondeterminism via bound checks or honest-majority voting.
  4. (4)
    Deployable PyTorch runtime and end-to-end system. We implement a PyTorch-compatible runtime that instruments models to compute and apply these error models in real time, integrate it into the full NAO pipeline, and demonstrate deployability in the open-model setting.
Outline of the paper. We first present our system architecture and protocol overview in Sec. 3, followed by detailed methodology of two error models in Sec. 4, the adversary model and attack experiments in Sec. 5. We introduce the optimistic dispute game in Sec. 6 and evaluate our system’s performance, security, and cost in Sec. 7. Related work is in Sec. 8 with discussion and conclusion in Sec. 9.
2. Background
2.1. Nondeterminism is the default (replication is brittle)
In production, floating-point execution order shifts with GPU count/type and library stacks (cuBLAS/cuDNN/ATen), drivers, and runtimes. Bitwise replay occurs only when all knobs match: a condition that is impractical at scale and in decentralized markets. Forcing determinism often sacrifices performance and hardware heterogeneity. We empirically verified that endpoint settings meant to reduce randomness do not guarantee identical outputs across invocations. Concretely, on the OpenAI API we issued the same prompt to the same model with temperature T=0T{=}0, nucleus parameter top-p=1p{=}1, number of candidates n=1n{=}1, and a fixed seed $s$. Despite controlling both input randomness and output randomness (sampler), we observed different token sequences on repeated calls; when requesting token-level log probabilities, the reported logprobs also differed across runs. Similar caveats apply to other hosted services: Amazon Bedrock’s documentation states responses can be made more deterministic but not fully deterministic. [^1] https://repost.aws/knowledge-center/bedrock-prompts-deterministic-responses These observations highlight a core constraint for verifiers: exact replay at the service boundary is brittle and costly, even before accounting for heterogeneous hardware.
2.2. Benign numerical drift is ubiquitous and regular
NAO takes a different path: instead of suppressing nondeterminism, we model the numerical variance caused by benign reorderings. We measure post-operator runtime error, i.e., the element-wise deviation between two valid executions of the same operator given the same inputs but on different (hardware, library) configurations. Using traced models, we ran the identical operator sequences on two GPU types (H100 and RTX6000) and recorded the intermediate operator outputs along the graph’s canonical topological order. Fig. 2 shows that per-op deviations are small, largely flat across depth, and feature only localized spikes. Per-operator error histograms (not shown) are sharply peaked with light tails. This regularity suggests an opportunity: tolerance-aware verification that accepts outputs within principled, op-specific acceptance regions without fixing a global execution order.
Figure 2. Mean empirical errors vs. normalized operator position (log-scale y-axis). The x-axis follows the canonical topological order. Profiles remain essentially flat with localized spikes, indicating limited room for attack.
Figure 3. NAO overview. Phase 0 (Model Setup): model owner commits to weights, graph, empirical error percentile thresholds, and per-operator trace schema. Phase 1 (Optimistic Execution): proposer posts a commitment for $(x,y)$. Phase 2 (Dispute Localization): an NN-way, Merkle-anchored dispute game recursively narrows disagreement to a single operator following a canonical partition policy. Phase 3 (Single-Operator Adjudication): either a sound IEEE-754 error-bound check or a committee vote against empirical thresholds decides the outcome, triggering payment or slashing.:
2.3. Sound theory alone is too loose (and exploitable)
Given the small empirical errors above, a natural baseline is to instantiate a compositional framework for theoretical IEEE-754 rounding bounds. For each operator with its input, we compute a worst-case, element-wise rounding envelope that is sound under standard floating-point error models. However, theory-only envelopes are far looser than what we observe in practice (comparison in Fig. 6). This slack enlarges the attack surface. With a traced model and per-op theoretical budgets in hand, an adversary can inject perturbations within these element-wise bounds after each op. Using projected-gradient steps on the task loss, they can steer the cumulative budget along sensitive directions to induce label flips or semantic drift (Fig. 1). On large models, such “theory-only” checks admit a small but non-trivial attack success rate (Table 2), demonstrating that worst-case soundness alone is insufficient for economically robust verification.
Bridge: closing the gap. The real problem therefore demands new ideas: eliminating the exploitable slack those bounds leave in real deployments. NAO closes this gap by calibrating empirical, per-model, per-operator error percentile thresholds across heterogeneous hardware and using them to guide dispute localization; leaves are then adjudicated cheaply via either certified theoretical checks or small committee votes against the stored empirical thresholds. This dual design embraces nondeterminism, removes replication burden, and makes verification practical and defensible. Later sections detail the mechanism.
In this section, we define the roles, assumptions, and the end-to-end protocol that makes NN results verifiable.
3. System architecture and protocol overview
System at a glance. NAO provides verifiable results for NN workloads by combining optimistic-rollup-style economics with tolerance-aware verification. Fig. 3 shows the four phases: in Phase 0 (Setup), the model owner commits Merkle roots of the weights and graph, and publishes calibrated per-operator empirical error percentile thresholds. In Phase 1 (Optimistic execution), a proposer runs the model to fulfill user’s request and the result finalizes after a short challenge window. If challenged, it runs a Merkle-anchored, threshold-guided interactive Phase 2 (dispute game), that recursively narrows disagreement from the entire model graph to a single PyTorch operator. Finally, Phase 3 (Single-operator adjudication) executes either (i) a sound, cheap but potentially permissive IEEE-754 theoretical-bound check, or (ii) a tighter but costlier committee vote against the empirical error percentile thresholds. Violations result in slashing.
3.1. Roles and assumptions
Roles. The system involves five roles:
$\bullet$ User: specifies model, submits inputs and requests results.
$\bullet$ Proposer: executes the model on heterogeneous accelerators and posts the results with a commitment. If the result is questioned, it participates in dispute games.
$\bullet$ Challenger: verifies result by re-executing the program. If its reproduced output deviates by the empirical error percentile threshold, it initiates and advances the disputes.
$\bullet$ On-chain contract: escrows payments, records commitments, enforces timeouts, and orchestrates the state machine.
$\bullet$ Committee: sampled participants that adjudicate the final single operator against empirical error percentile thresholds when the theoretical check is inconclusive.
Assumptions. We assume IEEE-754 compliance with round-to-nearest-even; fast-math is disabled; operator libraries may reorder reductions but must respect IEEE semantics. When invoked, the committee has an honest majority, and no trusted hardware is required.
3.2. End-to-end protocol lifecycle
In this section, we walk through the four phases: from commitment to adjudication, showing how disputes are localized and resolved in a high-level overview (Fig. 3).
Phase 0: Model setup. At deployment time, the model owner serializes the PyTorch model into an acyclic dataflow graph $G=(V,E)$ with a canonical topological order, where each node $v_{i}\in V$ denotes a tensor operator (e.g., MatMul, Softmax, GELU) and each edge $(u,v)\in E$ captures a data dependency. The model state is then committed on-chain by merkleizing the weight tensors to obtain root $r_{w}$ and, separately, merkleizing the graph topology (operator signatures) to obtain root $r_{g}$. The pair $(r_{w},r_{g})$ is recorded on the blockchain, enabling lightweight provenance proofs for any subgraph that arises during later dispute phases (Sec. 6.2). We also perform an offline calibration (Sec. 4.2) to quantify cross-hardware floating-point variability per operator, yielding empirical error percentile thresholds, which are also committed as a root $r_{e}$. These commitments enable verifiable subgraph extraction: any subgraph $S\subseteq G$ carries membership proofs into $r_{g}$ and $r_{w}$.
[…]
Phase 2: Dispute localization (interactive decomposition). If any verifier recomputes $y^{\prime}=G(x)$ and finds that the discrepancy between $y$ and $y^{\prime}$ exceeds the empirical error percentile threshold, they may initiate a dispute game and act as a challenger. The contract then freezes proposer’s collateral $\sigma_{P}$, starts a dispute game, and initializes round k=0k{=}0 with $S_{0}{:=}G$. Each round proceeds (details in Sec. 6.3):
$\bullet$ Partition (deterministic). The proposer partitions the current disputed subgraph $S_{k}$ into $N$ disjoint children $\{S_{k}^{j}\}(0\leq j<N)$ and posts for each child a compact commitment (subgraph inclusion Merkle proof and I/O hashes).
$\bullet$ Selection (first offending child). The challenger recomputes the same partitions locally and designates the first offending child $S_{k}^{j^{\star}}$ in topological order over $V$ guided by empirical error percentile thresholds, thereby implicitly agreeing on all earlier children and on the inputs to $S_{k}^{j^{\star}}$.
$\bullet$ Advance. The contract updates the dispute game state: sets $S_{k+1}\leftarrow S_{k}^{j^{\star}}$ and increments $k$, and enforces per-round timeouts. Failure to act within the timeout loses that round.
The game repeats until $|S_{k}|=1$, at which point the dispute has been localized to a single operator and Phase 3 is invoked. The number of rounds is $O(\log_{N}|V|)$.
Phase 3: Single-operator adjudication. At the leaf, both parties agree on the operator type $v^{*}$ and the inputs $a$. The challenger then tests for fraud using one of two methods: (i) a theoretical-bound check, where a verifiable VM computes element-wise FP error bounds (Sec. 4.1): If the proposer’s claimed outputs fall outside these bounds for any element, the contract slashes $\sigma_{P}$ and rewards the challenger. (ii) When the bound is inconclusive (the claim lies within the theoretical error bound), a committee check, where multiple independent runners execute $(v^{*},a)$ and vote using the pre-calibrated empirical error percentile thresholds: this test is much tighter but more costly.
4. Floating-point rounding error bounds and empirical thresholds for neural networks
In this section, we compute per-operator theoretical floating-point error bounds and calibrate empirical error percentile thresholds, enabling tolerance-aware per-operator verification across devices.
4.1. Computing theoretical error bound
This section derives first-order operator-local rounding-error bounds from the IEEE-754 model and shows how to compute them efficiently via FX-based co-execution templates.
[…]
Crucially, we do not propagate bounds across operators (network-level propagation is impractical for deep models); we only account for (i) error propagated within operator’s own sub-steps and (ii) fresh rounding errors from those steps.
We adopt the standard floating-point model as follows: for ${+,,×,/}\circ\in\{$$\times$,/\}, $\operatorname{fl}(x\circ y)=(x\circ y)(1+\delta)$ with $|\delta|\leq u$ ($u$ is unit roundoff). For library intrinsics we use vendor-stated maximum-ULP errors for CUDA math functions (Guide, 2013). For reductions composed of $k$ basic ops, we support both worst-case deterministic  (Higham, 2002; Goldberg, 1991) and relatively tighter probabilistic (Madry et al., 2018; Higham and Mary, 2019) bounds. A canonical example is sequential summation $s=\sum_{i=1}^{n}x_{i}$ (k=n1k=n-1):
$\text{deterministic: }\;|\widehat{s}-s|\leq\gamma_{n-1}\sum_{i=1}^{n}|x_{i}|,\quad\gamma_{k}\triangleq\frac{ku}{1-ku};$
$\text{probabilistic: }\;\Pr\!\Big(|\widehat{s}-s|\leq\widetilde{\gamma}_{n-1}(\lambda)\sum_{i=1}^{n}|x_{i}|\Big)\geq 1-2e^{-\frac{\lambda^{2}(1-u)^{2}}{2}},$
with $\widetilde{\gamma}_{k}(\lambda)\triangleq\exp\!\big(\lambda\sqrt{k}u+\frac{ku^{2}}{1-u}\big)-1$. We use $\lambda=4$, giving $\widetilde{\gamma}_{k}(4)\approx 4u\sqrt{k}$ at $\geq 99.93\%$ confidence; more details are shown in supplemental materials.
FX-based co-execution and operator templates. We extend PyTorch FX (PyTorch Contributors, 2025) interpreter to co-execute values and bounds. Each traced operator is lowered to a short sequence of primitives (adds/multiplications/reductions/intrinsics). For a primitive $f(x_{1},\dots,x_{m})$ with current inputs $\{x_{i}\}$ and already-accumulated intra-operator input errors $\{\varepsilon_{x_{i}}\}$, we use a first-order sensitivity envelope
[…]
and set $\varepsilon_{f}=\varepsilon^{\text{prop}}_{f}+\varepsilon^{\text{rnd}}_{f}$. Reductions (sum/mean/matmul/conv) use $\gamma_{k}$ or $\widetilde{\gamma}_{k}$ with the appropriate $k$ per output element for the fresh rounding term; pure data movement (e.g., view/cat/gather) contributes no FP error.
Minimal example: softmax. For a vector $x$ along dimension $d$, we compute the softmax operator $\hat{y}$ as:
$m=\max(x),\;z=x-m,\;e=\exp(z),\;S=\textstyle\sum\nolimits_{j}e_{j},\;\hat{y}_{i}=e_{i}/S.$
The per-step bounds (all elementwise) separate propagated vs. fresh rounding (first-order worst-case model):
$\varepsilon_{z}$ $\leq u\,(|x|+|m|),\qquad\varepsilon_{e}\leq|e|\,\varepsilon_{z}\;+2u\,|e|,$
$\varepsilon_{S}$ $\leq\widetilde{\gamma}_{n-1}\!\sum_{j}|e_{j}|+(\widetilde{\gamma}_{n-1}+1)\sum_{j}|\varepsilon_{e_{j}}|,$
$\varepsilon_{\hat{y}_{i}}$ $\leq\frac{\varepsilon_{e_{i}}}{|S|}\;+\;\frac{|e_{i}|\,\varepsilon_{S}}{|S|^{2}}\;+\;u\,|\hat{y}_{i}|.$
Here the derivative terms $|\partial\exp/\partial z|=|e|$ and $|\partial(e/S)/\partial(\cdot)|$ generate the propagated components, while $u|\cdot|$ terms are fresh rounding. Analogous templates cover other 35 common operators (LayerNorm, MatMul, Conv2d, etc.); We ignore the additional numerical errors introduced while computing these error bounds themselves.
Figure 4. Comparison between deterministic and probabilistic theoretical error bounds for different operator types in Qwen-8B and BERT-large. :
Putting it together. Our library of operator templates yields $\tau_{v}^{\text{theo}}$ cheaply during a single traced run. These per-operator bounds are later used for single-operator verification. Fig. 4 contrasts deterministic $\gamma_{k}$ with probabilistic $\widetilde{\gamma}_{k}$ across representative operators; the probabilistic bounds are markedly tighter, especially at large $k$. However, worst-case $\gamma_{k}$ and even $\widetilde{\gamma}_{k}$ remain conservative for real ML workloads where mixed signs induce cancellation; empirically, observed cross-device discrepancies are often orders of magnitude smaller. We therefore pre-calibrate per-operator error percentile profiles on representative inputs to construct empirical thresholds and use them alongside the theoretical bounds.
4.2. Calibrating empirical error percentile threshold
In this section, we calibrate device-agnostic absolute/relative error percentile profiles for each operator across inputs and hardware. We then validate their stability and inflate them into practical thresholds with a safety margin.
The dispute protocol narrows disagreements from the full computation graph to individual operators by steering a N-way partition search with empirically calibrated error percentile thresholds. Across inputs, intermediate tensor shapes may differ, so for feasible comparisons, we calibrate per-operator error distributions rather than using element-wise bounds. We describe our offline calibration in this section.
Cross-device calibration procedure. For a traced model $G=(V,E)$, we sample a representative dataset $\mathcal{D}=\{x_{1},\ldots,\allowbreak x_{m}\}$ and evaluate it on multiple GPU devices $\mathcal{H}=\{h_{1},\ldots,h_{k}\}$. For operator $v_{i}\in V$ and input $x_{s}$, let $y_{i}^{(h_{j},x_{s})}\in\mathbb{R}^{d_{i}}$ denote the output tensor of $v_{i}$ on device $h_{j}$, where $d_{i}$ is the tensor dimension after the tensor being flattened.
For each device pair $(h_{j},h_{k})$, we form element-wise absolute and relative errors:
(1) $\mathbf{e}_{\mathrm{abs}}^{(i,j,k,s)}$ $=|y_{i}^{(h_{j},x_{s})}-y_{i}^{(h_{k},x_{s})}|\in\mathbb{R}^{d_{i}},$
(2) $\mathbf{e}_{\mathrm{rel}}^{(i,j,k,s)}$ $=\frac{|y_{i}^{(h_{j},x_{s})}-y_{i}^{(h_{k},x_{s})}|}{|y_{i}^{(h_{j},x_{s})}|+\epsilon}\in\mathbb{R}^{d_{i}},$
with a small $\epsilon>0$ to avoid division by zero.
Percentile profiles. For each $(i,j,k,s)$, we compute the percentile-value vector over tensor elements at percentiles $\allowbreak p\allowbreak\in\allowbreak P\allowbreak=\allowbreak\{\allowbreak 0\allowbreak,\allowbreak 1\allowbreak,\allowbreak 5\allowbreak,\allowbreak 1\allowbreak 0\allowbreak,\allowbreak 1\allowbreak 5\allowbreak,\allowbreak\ldots\allowbreak,\allowbreak 8\allowbreak 5\allowbreak,\allowbreak 9\allowbreak 0\allowbreak,\allowbreak 9\allowbreak 5\allowbreak,\allowbreak 9\allowbreak 9\allowbreak,\allowbreak 1\allowbreak 0\allowbreak 0\allowbreak\}\allowbreak$:
(3) $\mathcal{P}_{\mathrm{abs}}^{(i,j,k,s)}(p)$ $=\text{Percentile}_{p}(\mathbf{e}_{abs}^{(i,j,k,s)}),$
(4) $\mathcal{P}_{\mathrm{rel}}^{(i,j,k,s)}(p)$ $=\text{Percentile}_{p}(\mathbf{e}_{rel}^{(i,j,k,s)}),$
where $\mathcal{P}(\cdot)$ denotes the vector of percentile profiles (error distribution) indexed by $p\in P$, and is computed over all $d_{i}$ elements of the error tensor. We obtain conservative, device- and input-agnostic operator-wise profiles by a max-envelope across $j,k,s$:
(5) $\mathcal{P}_{\mathrm{abs}}^{(i)}(p)$ $=\max_{j,k}\left\{\{\mathcal{P}_{\mathrm{abs}}^{(i,j,k,s)}(p)\}_{s=1}^{m}\right\},$
(6) $\mathcal{P}_{\mathrm{rel}}^{(i)}(p)$ $=\max_{j,k}\left\{\{\mathcal{P}_{\mathrm{rel}}^{(i,j,k,s)}(p)\}_{s=1}^{m}\right\}.$
Table 1. Stability metrics at selected percentiles (p30, p50, p70) for three models. All values are normalized by median.
Model $p$ SupNorm Jackknife TailAdj RollSD
@50 @90 @50 @90 @50 @90 @50 @90
Qwen 30 0.00 0.03 0.00 0.02 0.00 0.03 0.05 0.11
50 0.00 0.02 0.00 0.02 0.00 0.03 0.05 0.11
70 0.00 0.02 0.00 0.02 0.01 0.03 0.05 0.10
BERT 30 0.00 0.04 0.00 0.00 0.00 0.03 0.03 0.10
50 0.00 0.04 0.00 0.00 0.00 0.03 0.03 0.09
70 0.00 0.04 0.00 0.01 0.00 0.02 0.03 0.08
ResNet 30 0.00 0.02 0.00 0.00 0.00 0.03 0.00 0.10
50 0.00 0.04 0.00 0.01 0.00 0.03 0.00 0.09
70 0.00 0.05 0.00 0.01 0.00 0.03 0.00 0.10
Statistical validation. We sampled 50 input data for each model to collect the percentile profiles of each operator. We assess per-operator stability with four metrics: SupNorm (drift), Jackknife (leave-one-out influence), TailAdj (recent tail change), and RollSD computed over running medians (window $W=10$) and summarized at the 50th/90th percentiles across operators, normalized by each metric’s median (Table 1; details in Sec. B). Across three models and $p\!\in\!\{30,50,70\}$, central tendencies are $\approx 0$, and 90th-percentile bounds are tight (SupNorm $\leq 0.05$, Jackknife $\leq 0.02$, TailAdj $\leq 0.03$); RollSD is modestly higher (up to $0.11$). These indicate near-stationary operator estimates and support the reliability of our empirical error profiles. The metric details are in supplementary materials.
Threshold construction. To account for calibration uncertainty and provide a safety margin, we apply a multiplicative factor $\alpha=3$ to every percentile value for each $v_{i}\in V$:
(7) $\tau_{\mathrm{abs}}^{(i)}(p)=\alpha\cdot\mathcal{P}_{\mathrm{abs}}^{(i)}(p),\qquad\tau_{\mathrm{rel}}^{(i)}(p)=\alpha\cdot\mathcal{P}_{\mathrm{rel}}^{(i)}(p).$
The set $\{\tau_{\mathrm{abs}}^{(i)},\tau_{\mathrm{rel}}^{(i)}\}_{i=1}^{|V|}$ is bundled with the model commitment $(r_{w},r_{g})$ and is included in the on-chain commitment.
Taken together, theoretical bounds provide sound, portable bounds, while empirical profiles yield tight, model-specific guidance. This combination enables well-posed, tolerance-aware verification that respects IEEE-754 nondeterminism without sacrificing production performance.
5. Threat model and attack design
This section formalizes the adversary and our security objectives, and states that in NAO the only attack surface is the two error bounds. We then describe the attacks used to probe these interfaces under their respective conditions and, finally, demonstrate the robustness of the error bounds.
5.1. Adversary and objectives
Adversary capabilities. We consider only the proposer adversarial; all other parties follow the protocol. The adversary is white-box: they know model commitment, input, canonical topological order, empirical error percentile thresholds, and the dispute-partition policy. The adversary may alter intermediate tensors during the computation but cannot change weights or graph topology without violating Merkle commitments, nor influence committee sampling or cryptography.
Security objectives. We evaluate: (i) Soundness: deviations that beyond committed bounds are detected (by empirical thresholds or theoretical bounds); (ii) Non-accumulation: admissible per-operator deviations do not compose into large task-level error; (iii) Low false positives: honest executions rarely trigger disputes under empirical thresholds.
5.2. Attack surface and empirical headroom
The attack surface is numerical: at selected $v$, the proposer injects additive perturbations $h_{v}\leftarrow h_{v}+\Delta_{v}$ to change the final decision while remaining within challenger’s admissible sets. To quantify headroom, we trace mean cross-device error per operator and plot it against normalized depth (Fig. 2). Across BERT-large, Qwen-8B, and ResNet-152, errors are essentially flat with localized spikes: we observe no systematic accumulation of error as depth increases. Typical magnitudes are $10^{-6}$-$10^{-5}$. Thus, flipping outputs while staying admissible requires coordinating many sub-micro deviations; over-allocation will be isolated by the dispute game.
5.3. Evasion strategies
In this section, we define two evasion strategies aligned with search-time and leaf-time verification checks: the adversary chooses $\{\Delta_{v}\}$ to change the final decision while remaining within the verifier’s admissible sets.
Empirical-threshold evasion (search time). For operator $v_{i}$, calibration commits percentile-cap pairs $\{(p_{k},\tau_{\mathrm{abs}}^{(i)}(p_{k}))\}_{k=1}^{K}$ and relative caps $\tau_{\mathrm{rel}}^{(i)}(\cdot)$. Define the nondecreasing cap curve $C_{i}:[0,1]\!\to\!\mathbb{R}_{\geq 0}$ by linear interpolation through $(0,0)$, $(p_{k},\tau_{\mathrm{abs}}^{(i)}(p_{k}))$, and $(1,\tau_{\mathrm{abs}}^{(i)}(1))$. The feasible set is
(8) $\mathcal{F}^{\text{emp}}_{v_{i}}\triangleq\Big\{\Delta:\ Q_{|\,\Delta\,|}(r)\leq C_{i}(r),\ \forall r\in[0,1]\Big\},$
where $Q_{|\,\Delta\,|}$ is the empirical quantile function of $|\Delta|$. Evasion requires the observed proposer–challenger discrepancy at each inspected operator to stay within $\mathcal{F}^{\text{emp}}_{v_{i}}$ for given $p_{k}$}.
Theoretical-bound evasion (leaf check). At the leaf, the challenger either runs the theoretical element-wise IEEE-754 bound or invokes a small committee vote. Here we focus on the attack against the theoretical bound check (latter is the empirical threshold check). For operator $v^{\star}$ with bound $\tau^{\text{theo}}_{v^{\star}}$, the feasible set is
(9) $\mathcal{F}^{\text{theo}}_{v^{\star}}\triangleq\big\{\Delta:\ |\Delta|\leq\tau^{\text{theo}}_{v^{\star}}\ \text{element-wise}\big\}.$
Because these worst-case bounds are typically looser than empirical error thresholds, the protocol falls back to a committee vote against empirical thresholds when the theoretical check is inconclusive.
5.4. Optimization-based attack method
We implement projected-gradient attack (Madry et al., 2018) over $\{\Delta_{v}\}$ to flip the predicted class/token while staying within a chosen admissible set.
Attack objective. Let $z\in\mathbb{R}^{C}$ be the logits at model’s output, and we maximize the logit margin:
(10) $L_{\text{margin}}(z)\;=\;z_{c_{2}}-z_{c_{1}},$
where $c_{1}=\arg\max_{c}z_{c}$ and $c_{2}$ is the attack’s target class. It succeeds once $L_{\text{margin}}>0$.
PGD updates and projections. At iteration $t$, compute $g_{v}^{t}=\nabla_{\Delta_{v}}L_{\text{margin}}$ through the traced graph, apply Adam with $(\beta_{1},\beta_{2},\varepsilon)=(0.9,0.999,10^{-8})$ and a per-operator stepsize $\eta_{v}$ (default: $1/4$ median of its error bound), to get a tentative $\widetilde{\Delta}_{v}^{\,t+1}$. We then project onto the relevant feasible set:
(i) Projection under theoretical bounds. This projection is simply element-wise clipping ($\tau_{v}$ is computed at runtime using this operator’s input):
(11) $\Delta_{v}^{t+1}=\operatorname{clip}\!\left(\widetilde{\Delta}_{v}^{\,t+1},-\tau_{v},\tau_{v}\right).$
(ii) Projection under empirical thresholds. Let $a=\bigl|\mathrm{vec}(\widetilde{\Delta}_{v}^{\,t+1})\bigr|\in\mathbb{R}^{n}$, $s=\mathrm{sign}(\mathrm{vec}(\widetilde{\Delta}_{v}^{\,t+1}))$, and let $\sigma$ sort $a$ increasingly. For ranks $r_{k}=(k-\tfrac{1}{2})/n$, define caps $c_{k}=C_{v}(r_{k})$ and enforce monotonicity via $c_{k}\leftarrow\max(c_{k},c_{k-1})$. Clip order statistics $a^{\star}_{\sigma(k)}=\min\{a_{\sigma(k)},c_{k}\}$, then restore sign/shape:
(12) $\Delta_{v}^{t+1}=\mathrm{reshape}\;\!\big(s\odot a^{\star}\big).$
This realizes the projection onto $\mathcal{F}^{\text{emp}}_{v}$. The dominant cost is the sort: $O(n\log n)$ per operator; all steps are vectorized.
Table 2. Bucketed attack outcomes under threshold scaling. Each cell reports ASR (%) and mean $\Delta m_{\text{fail}}$ ($\delta_{\text{fail}}$). Buckets group targets by their logit margin percentile. Scale $\alpha$ multiplies the error bounds (larger is looser). False Positive (%) is the honest-run dispute rate at the same $\alpha$. (For theoretical bound attack, d is deterministic bound; p is probabilistic bound)
Model Bound check Scale $\alpha$ 0–20% 20–40% 40–60% 60–80% 80–100% False Positive (%)
ASR (%) $\overline{\Delta m}_{\text{fail}}$($\overline{\delta}_{\text{fail}}$) ASR (%) $\overline{\Delta m}_{\text{fail}}$($\overline{\delta}_{\text{fail}}$) ASR (%) $\overline{\Delta m}_{\text{fail}}$($\overline{\delta}_{\text{fail}}$) ASR (%) $\overline{\Delta m}_{\text{fail}}$($\overline{\delta}_{\text{fail}}$) ASR (%) $\overline{\Delta m}_{\text{fail}}$($\overline{\delta}_{\text{fail}}$)
BERT-large Empirical $\times$1 0.0 0.02(0.3%) 0.0 0.02(0.2%) 0.0 0.02(0.2%) 0.0 0.02(0.2%) 0.0 0.02(0.2%) 0
$\times$2 0.0 0.04(0.5%) 0.0 0.04(0.4%) 0.0 0.04(0.4%) 0.0 0.04(0.4%) 0.0 0.03(0.3%) 0
$\times$3 0.0 0.06(0.8%) 0.0 0.06(0.6%) 0.0 0.06(0.6%) 0.0 0.06(0.6%) 0.0 0.05(0.5%) 0
Theo (d/p) $\times$1(d) 58.6 8.34(94%) 32.6 8.32(88%) 26.4 8.56(88%) 20.8 8.73(87%) 14.6 8.73(84%) -
$\times$1(p) 0.0 0.71(8.9%) 0.0 0.62(6.9%) 0.0 0.64(6.9%) 0.0 0.61(6.3%) 0.0 0.56(5.6%) -
$\times$0.5(p) 0.0 0.32(4.1%) 0.0 0.28(3.1%) 0.0 0.28(3.0%) 0.0 0.27(2.8%) 0.0 0.24(2.4%) -
ResNet-152 Empirical $\times$1 0.0 0.31(4.0%) 0.0 0.30(3.8%) 0.0 0.30(3.7%) 0.0 0.30(3.6%) 0.0 0.30(3.6%) 0
$\times$2 0.0 0.62(8.0%) 0.0 0.60(7.4%) 0.0 0.59(7.2%) 0.0 0.60(7.1%) 0.0 0.60(7.0%) 0
$\times$3 0.0 0.94(12%) 0.0 0.91(11%) 0.0 0.89(11%) 0.0 0.89(11%) 0.0 0.90(10%) 0
Theo (d/p) $\times$1(d) 0.0 2.67(34%) 0.0 2.54(31%) 0.0 2.50(30%) 0.0 2.49(29%) 0.0 2.46(29%) -
$\times$1(p) 0.0 0.50(6.7%) 0.0 0.48(6.1%) 0.0 0.48(6.0%) 0.0 0.48(5.9%) 0.0 0.48(5.7%) -
$\times$0.5(p) 0.0 0.25(3.3%) 0.0 0.24(3.0%) 0.0 0.24(3.0%) 0.0 0.24(2.9%) 0.0 0.24(2.8%) -
Qwen-8B Empirical $\times$1 0.0 0.77(4.2%) 0.0 0.79(3.8%) 0.0 0.81(3.6%) 0.0 0.82(3.5%) 0.0 0.87(3.4%) 0
$\times$2 0.0 1.45(7.9%) 0.0 1.46(7.0%) 0.0 1.51(6.7%) 0.0 1.54(6.5%) 0.0 1.60(6.3%) 0
$\times$3 0.0 2.05(11%) 0.0 2.08(9.8%) 0.0 2.14(9.5%) 0.0 2.18(9.2%) 0.0 2.26(8.9%) 0
Theo (d/p) $\times$1(d) 12.6 18.9(98%) 9.2 19.4(96%) 6.4 19.7(95%) 9.2 20.5(94%) 8.2 21.2(89%) -
$\times$1(p) 2.4 7.31(40%) 2.4 7.46(36%) 1 7.76(35%) 1.2 8.05(34%) 0.6 8.35(33%) -
$\times$0.5(p) 0.6 5.87(31%) 0.2 5.89(28%) 0.8 6.02(27%) 0.2 6.13(26%) 0.6 6.36(25%) -
5.5. Attack evaluation
In this section, we measure attack success, progress, and false positives across models under threshold scaling. Experiments are conducted on ResNet-152, BERT-large, and Qwen3-8B.
Datasets and hyperparameters. For BERT-large (Devlin et al., 2019) we use DBpedia (Auer et al., 2007); for Qwen3-8B (Team, 2025) we use C4 (Raffel et al., 2019), feeding the first 10% of each sequence and targeting the next token; for ResNet-152 (He et al., 2016) we use ImageNet (Deng et al., 2009) validation set. PGD/Adam and projection operators follow Sec. 5.4. Early stopping triggers when the last 10 updates satisfy $|m_{t}-m_{t-1}|<10^{-3}|m_{0}|$ and $|m_{t}|<10^{-3}|m_{0}|$ ($m_{t}$ is the logit margin at time step $t$). We use 500 samples (2500 attack targets) for the theoretical-bound attack and 250 (1250 attack targets) for the empirical-threshold attack to balance statistical precision and computational cost.
Metrics and bucketing. We report (i) attack success rate (ASR), the fraction of attacks that successfully flip the model’s prediction to the target class while staying under the chosen bound, and (ii) the mean logit margin change on attack failed runs, $\overline{\Delta m}_{\text{fail}}$, measuring how close failed attacks came to success, and $\overline{\delta}_{\text{fail}}$ measures the normalized margin change to the target margin. For example, let $m_{0}=z_{c_{1}}-z_{c_{2}}>0$ be the logit margin between the original predicted token $c_{1}$ and the target token $c_{2}$ before attack begins. After the attack process, the new logit gap is $m^{\prime}=z^{\prime}_{c_{1}}-z^{\prime}_{c_{2}}$. Then, $\Delta m=m_{0}-m^{\prime}$ and $\delta=\frac{\Delta m}{m_{0}}$. To study the perturbation for different levels of token / classes, for each model input, we bucket and sample examples by the target’s logit margin in the original prediction (lower means smaller margins): $0,20\%],[20\%,40\%],[40\%,60\%],[60\%,80\%],[80\%,100\%]$. We randomly sampled one class from each bucket to attack. We also scale bounds by a factor $\alpha$ to probe sensitivity: $\alpha>1$ loosens empirical error percentile thresholds; $\alpha<1$ tightens theoretical error bounds (diagnostic only). For theoretical error bounds we show d = deterministic theoretical bounds and p = probabilistic theoretical bounds (Sec. 4). This comparison also shows the reason for selecting probabilistic bounds in practice. False Positive (FP) is measured by running the full pipeline on honest executions across 1k inputs sampled from the dataset and reporting the rate at which the pipeline raises a dispute against honest computations.
Results and analysis. Table 2 summarizes the attack outcomes. We can conclude the following:
(1) Empirical-threshold is robust across models and all target attack settings. With empirical error percentile thresholds, ASR keeps $0.0\%$ for all three models even when bounds are loosened by $3\times$ ($\alpha=3$). The normalized progress on failed attack is minimal, with $\overline{\delta}_{\text{fail}}<1\%$ for BERT-large and $\overline{\delta}_{\text{fail}}<10\%$ for ResNet-152 and Qwen3-8B. It demonstrates the effectiveness of our empirical threshold verification.
(2) Theoretical bound check leaves a small attack window for LLMs. Under the theoretical element-wise bounds at the leaf, BERT-large and ResNet-152 still maintain $0.0\%$ ASR, while Qwen3-8B exhibits low yet non-zero vulnerability: up to $2.4\%$ ASR at $\alpha=1$ and $0.6\%$ ASR at $\alpha=0.5$. The mean progress on failed attacks is significant, with $\overline{\delta}_{\text{fail}}$ up to $40\%$ at $\alpha=1$. These cases align with the intuition that larger models (LLMs) can be sensitive to small logit perturbations; worst-case IEEE-754 theoretical bounds are loose and may not always rule the malicious perturbations out. It also justifies our choice of probabilistic theoretical bounds over deterministic ones, as the latter leaves a much larger attack window. The comparison between (1) and (2) also highlights the advantage of empirical thresholds over theoretical worst-case bounds in practice for a meticulous attacker.
(3) No false positives with empirical bounds. Across all models and scaling factors, the full dispute pipeline with empirical bounds never raises a dispute against honest nodes, indicating that our empirical bounds are not only robust against attacks but also practical for real-world deployment.
Figure 5. Normalized margin change on failed attacks ($\alpha{=}1$). Empirical thresholds yield near-zero progress across models; theoretical bounds(p) show heavier tails for the LLM. Numbers above boxes show sample sizes.:
(4) Margin-change distributions corroborate tightness. Fig. 5 presents a boxplot of the normalized margin change for the failed cases after the attack. The empirical threshold attacks are all tightly concentrated at  0.05, indicating that admissible perturbations yield almost no progress towards flipping the decision. In contrast, the theoretical bound attack for Qwen-8B exhibits a pronounced step-up, indicating that some perturbations can make significant progress towards flipping its decision under the theoretical bounds.
Key insights. Empirical error percentile thresholds keep disputes localized and robust without determinism, while theoretical error bounds give a low-cost, portable fallback at the leaf. The tradeoff is model-dependent: for LLMs, worst-case IEEE-754 bounds can be permissive, motivating a committee vote when tighter adjudication is desired.
6. NAO: Optimistic operator-level dispute protocol
In this section, we adopt an optimistic design in which a public L1 blockchain serves as arbiter. A proposer commits to the model, inputs, and outputs; if no challenger disputes within a window, the result finalizes cheaply on L1. Upon dispute, an interactive localization game narrows disagreement to a single operator guided by the error tolerance, after which L1 performs a constant-cost check; the losing party is slashed.
6.1. Background and design choices.
Prior verification systems, including refereed delegation of computation, which guarantees correctness by querying multiple untrusted servers and resolving disagreements via a referee (Canetti et al., 2013, 2011), and interactive proofs (Goldwasser et al., 2008; Kalai et al., 2022). Rollup-style verification games (e.g., Truebit, Arbitrum, OP Stack (Teutsch and Reitwießner, 2024; Kalodner et al., 2018; Optimism, 2025b, a)) instantiate this over CPU-style VMs via instruction-level bisection. NN on heterogeneous accelerators violates both: GPU kernels cannot be stepped by existing VMs and floating-point non-associativity plus parallel scheduling introduce cross-hardware variability (PyTorch Contributors, 2024; Higham, 2002).
NAO preserves the optimistic, interactive structure but replaces the units of dispute and the base-case check: (i) we perform N-way partition on a Merkle-anchored, operator-granular NN graph rather than a VM trace; (ii) we guide partition selection via tolerance-aware error thresholds calibrated per operator; and (iii) at the leaf we perform either a sound IEEE-754 theoretical-bound check or (when bounds are permissive) a committee vote against empirical error percentile thresholds. This design is new in (a) moving from instruction-level to operator-level partition, (b) defining a deterministic selection rule under non-deterministic floats, and (c) providing a constant-cost L1 adjudication path that respects IEEE semantics without requiring a GPU VM.
6.2. Operator-granular graph representation and verifiable subgraph extraction
In this section, we trace models to an operator-level dataflow graph and show how to extract, commit to, and verify Merkle-anchored subgraphs.
Cut sets and subgraph extraction. Given a candidate subgraph $S$ (a contiguous slice of operators in the execution order), we compute its live-in and live-out frontiers by a linear scan:
(13) $\mathrm{In}(S)\;$ $=\;\{\,u\notin S\mid\exists\,v\in S:(u\to v)\,\},$
(14) $\mathrm{Out}(S)\;$ $=\;\{\,w\notin S\mid\exists\,v\in S:(v\to w)\,\}.$
We materialize $S$ as an fx.GraphModule by cloning operators in $S$, replacing references to $\mathrm{In}(S)$ with placeholders, and emitting $\mathrm{Out}(S)$ as graph outputs. Parameters are reused by reference; operator attributes and tensor metadata are preserved.
Merkle commitments to weights, graph, and interfaces. We commit separately to weights and graph structure via Merkle trees whose roots are published on chain. We define $\mathsf{H}(\cdot)$ as the SHA-256 hash function.
Weights. We sort state_dict keys lexicographically and hash each parameter tensor:
$\mathcal{H}_{w}(k)=\mathsf{H}\big(\textsf{canon}(\texttt{state\_dict}[k])\big),$
where $\textsf{canon}(\cdot)$ serializes raw tensor bytes, dtype, shape, and stride, forming the weight Merkle tree $\mathcal{T}_{w}$ with root $r_{w}$.
Graph structure. For each operator $n$, the signature:
$\sigma(n)=\textsf{canon}\!\big(\mathrm{name}(n),\mathrm{op}(n),\mathrm{target}(n),\mathrm{args}(n),\mathrm{kwargs}(n)\big),$
captures kind, attributes, and topology (edges are implied by arguments). The leaf hash is $\mathcal{H}_{g}(n)=\mathsf{H}\big(\sigma(n)\big)$, forming the structure tree $\mathcal{T}_{g}$ with root $r_{g}$. This construction enables logarithmic-depth proofs, making on-chain verification feasible even for large models.
To bind interfaces, we generate the commitment for inputs and outputs of $S$:
$h_{D}=\mathsf{H}\!\big(\big\|_{z\in D(S)}\mathsf{H}(\textsf{canon}(z))\big),D\in\{\mathrm{In},\mathrm{Out}\}.$
Model configuration (percentile grid $P$ and thresholds $\tau$) is part of the commitment and cannot change mid-dispute.
Subgraph record and verification. For each extracted $S$, proposer publishes (i) its start/end indices in the topological order, (ii) $h_{\text{In}}$ and $h_{\text{Out}}$, and (iii) Merkle inclusion proofs for all leaves referenced in $S$ in $\mathcal{T}_{w}$ and $\mathcal{T}_{g}$. A challenger can verify by (1) recomputing $r_{w}$ from revealed leaves using $\mathcal{T}_{w}$ proofs, (2) checking membership of $S$’s operator signatures in $\mathcal{T}_{g}$, and (3) recomputing $h_{\text{In}},h_{\text{Out}}$ from observed tensors to check the integrity. Any tampering breaks a proof.
6.3. Threshold-guided dispute localization
In this section, we show how to deterministically select the offending model slice by comparing error-percentile profiles to calibrated per-operator thresholds.
Why naive selection fails. Instruction/time-step bisection is undefined for GPUs; across inputs, intermediate tensor shapes may differ: without a shape-aware, deterministic selection rule, two honest parties may disagree about which child is “offending”. Thus, we adopt a tolerance-aware selection rule based on empirical error percentile thresholds.
Tolerance-aware selection rule. Starting with $S_{0}\!=\!G$, proposer $P$ partitions $S_{k}$ into $N$ disjoint, contiguous slices $\{S_{k}^{0},\ldots,S_{k}^{N-1}\}$. For each slice $S_{k}^{j}$ and live-out operator $v_{i}\in\mathrm{Out}(S_{k}^{j})$, challenger $C$ re-executes $S_{k}^{j}$ and forms element-wise absolute and relative errors between proposer and its local outputs (flattened to 1-D):
$\mathbf{e}_{\mathrm{abs}}^{(i)}=\text{flatten}(|y_{i}^{(P)}-y_{i}^{(C)}|),\;\mathbf{e}_{\mathrm{rel}}^{(i)}=\text{flatten}(\frac{|y_{i}^{(P)}-y_{i}^{(C)}|}{|y_{i}^{(C)}|+\epsilon}),$
and computes the percentile profile over $\allowbreak p\allowbreak\in\allowbreak P\allowbreak=\allowbreak\{\allowbreak 0\allowbreak,\allowbreak 1\allowbreak,\allowbreak 5\allowbreak,\allowbreak 1\allowbreak 0\allowbreak,\allowbreak 1\allowbreak 5\allowbreak,\allowbreak\ldots\allowbreak,\allowbreak 9\allowbreak 5\allowbreak,\allowbreak 9\allowbreak 9\allowbreak,\allowbreak 1\allowbreak 0\allowbreak 0\allowbreak\}\allowbreak$:
$p_{\mathrm{abs/rel}}^{(i)}(p)$ $=\mathrm{Percentile}_{p}(\mathbf{e}_{\mathrm{abs/rel}}^{(i)})\,.$
For each $v_{i}$, the challenger compares the observed error percentiles against the calibrated thresholds:
(15) $p_{i}^{(max)}$ $=\max_{p\in P}\left\{\frac{p_{abs}^{(i)}(p)}{\tau_{abs}^{(i)}(p)},\frac{p_{rel}^{(i)}(p)}{\tau_{rel}^{(i)}(p)}\right\}.$
If any $p_{i}^{(max)}>1$, the discrepancy lies within or upstream of $S_{k}^{j}$. Challenger continues the search before $S_{k}^{j}$. $N$ way partition yields depth $O(\log_{N}n)$. The procedure stops when $|S_{k}|=1$, isolating a single disagreeing operator $v^{*}$ whose inputs and attributes are already fixed by prior commitments.
6.4. Single-operator verification
Routing policy. At the single operator $v^{*}$ (inputs and attributes already committed), the challenger chooses one of two verification paths. Routing policy: off-chain the challenger recomputes a reference $y_{v^{*}}^{\mathrm{ref}}$ and compares the elementwise error $|y_{v^{*}}^{P}-y_{v^{*}}^{\mathrm{ref}}|$ to the theoretical cap $\tau^{\text{theo}}_{v^{*}}$. If any element exceeds $\tau^{\text{theo}}_{v^{*}}$, it invokes path (i); otherwise, it invokes path (ii) to apply the tighter empirical thresholds.
(i) Theoretical-bound verification. The challenger requests the proposer to submit a verifiable proof (through opVM/zkVM) that, given the committed inputs and operator metadata: (i) a canonical reference execution computes $y_{v^{*}}^{\mathrm{ref}}$ under IEEE-754 FP32 semantics; (ii) the operator-wise theoretical bound $\tau^{\text{theo}}_{v^{*}}$ is computed; (iii) element-wise, $|y_{v^{*}}^{P}-y_{v^{*}}^{\mathrm{ref}}|\leq\tau^{\text{theo}}_{v^{*}}$. The on-chain verifier checks the single proof. If it passes the verification, the computation is accepted; otherwise, the proposer loses the dispute.
(ii) Committee vote with empirical thresholds. Each member in the committee re-executes that operator to get $y_{v^{*}}^{\mathrm{ref}}$, computes element-wise absolute/relative errors versus $y_{v^{*}}^{P}$, forms the error percentile profile, and compare against the calibrated error percentile thresholds $\{\tau^{(*)}_{\mathrm{abs}},\tau^{(*)}_{\mathrm{rel}}\}$ for $v^{*}$. If the majority of committee members vote that the proposer’s output is within the thresholds, it is accepted; otherwise, the proposer loses the dispute.
6.5. Economic soundness and incentives.
While the dispute game ensures correctness at the technical level, it can be combined with established economic mechanisms that allocate rewards and penalties to achieve incentive compatibility for proposers, challengers, and committee members, while also discouraging collusion. Such mechanisms have been extensively developed in other accountability frameworks for off-chain execution, including incentive-secure proof-of-learning protocols (Zhao et al., 2025), sampling-and-penalty schemes (Zhang et al., 2024), and watchtower-style networks (Sheng et al., 2024). We directly use these mature mechanisms to support our protocol: honest execution remains individually rational, voluntary challenges are properly incentivized, and committee participation under randomized audits is sustainable.
7. Implementation and evaluation
We implement NAO as (i) a PyTorch runtime that instruments graphs, computes bounds, splits/executes subgraphs, proceeds dispute games, and emits/verifies commitments, and (ii) a set of contracts on Ethereum Holesky orchestrating commitments, dispute state, per-round bonds. The runtime adds no custom kernels and runs unmodified models.
7.1. Experimental setup
Hardware and software configuration. Empirical error percentile threshold calibration uses four GPUs (NVIDIA RTX 4090, A100, H100, RTX 6000). We use PyTorch with FP32 forwards and FP64 for bound arithmetic. We set: fixed RNG seeds, fast-math disabled, CUDA/cuDNN library determinism, TF32/benchmark disabled, deterministic cuBLAS via CUBLAS_WORKSPACE_CONFIG. Contracts run on Ethereum Holesky testnet; we report on-chain costs in k gas. We ignore network latency during interactions.
Models and datasets. We evaluate ResNet-152 (He et al., 2016) on ImageNet (Deng et al., 2009), BERT-large (Devlin et al., 2019) on DBpedia (Auer et al., 2007), Qwen3-8B (Team, 2025) on C4 (Raffel et al., 2019) (next-token prediction), and Stable Diffusion v1-5 (Rombach et al., 2022). PyTorch graph sizes range from $1k$ to $5k$ operators.
Evaluation metrics. (i) Avg. dispute rounds: average number of rounds required to reach the leaf; (ii) Avg. dispute time: average off-chain dispute time (including both proposer and challenger with disk I/O); (iii) Avg. Merkle checks: average number of weight/topology proof verifications; (iv) per-round substep time during the dispute game (proposer partition vs. challenger selection). For this metric, we inject perturbations separately into eight operators in the model and record per-round substep time to compute the variance; (v) DCR: the challenger’s total FLOPs to reach the leaf during the dispute game; and (vi) Cost Ratio $\mathrm{DCR}/$model forward FLOPs, for normalized comparison.
7.2. System microbenchmarks and error analysis
Figure 6. Comparison between empirical and theoretical error bounds for different operator types in ResNet 152, Qwen-8B and BERT-large models.
Figure 7. Microbenchmarks on BERT-large dispute game: vary split size NN and dispute round index. Left: avg dispute rounds decrease roughly as $O(\log_{N}|V|)$. Mid-left: avg dispute time drops sharply from small $N$ and plateaus for $N\!\geq\!8$. Mid-right: avg Merkle proof checks for weights/graph fall as NN grows. Right: per-round substep time (proposer partition vs. challenger re-execute plus selection) measured across eight perturbed operators through the model.:
Error-bound statistics and relative tightness. In Fig. 6 we collect element-wise empirical and theoretical errors across all operators for three models to compare distributions:
$\bullet$ BERT-large. Mean empirical errors are extremely tight: $80\%$ of operators admit at most $10^{-6}$. The theoretical errors are much looser, with a mode at $10^{-3}$ covering $69\%$ of operators.
$\bullet$ Qwen3-8B. The empirical errors concentrate in the $10^{-5}$$10^{-6}$ bins, whereas theoretical errors are dominated by $10^{-2}$$10^{-3}$, again yielding a $10^{2}$$10^{3}$ gap.
$\bullet$ ResNet-152. Both empirical and theoretical errors are around $10^{-5}$, but the latter shows a longer permissive tail.
Implications for verification. First, search-time checks are one–three orders of magnitude tighter than worst-case theory for transformers; evasion would require perturbations that are consistently sub-micro across many operators, aligning with the non-accumulation results in Sec. 5. Second, leaf-only verification under theoretical bounds leaves more slack: especially for LLM stacks, motivating the committee option at the leaf when tighter adjudication is desired. We subsequently confirm these implications in the attack experiments, where empirical error percentile thresholds yield $0\%$ ASR across models while worst-case bounds allow a small attack window (Sec. 5.5).
Dispute game statistics and scaling with $N$. Fig. 7 varies the partition size $N$. Avg. Dispute Rounds: increasing $N$ reduces rounds from $\sim\!11$ at N=2N{=}2 to $\sim\!3$ at $N{\geq}12$. Avg. Dispute Time: dispute time drops sharply from N=2N{=}2 to $N{\approx}6$ and then flattens; the first round dominates latency because it handles the largest subgraph. On-chain work: Merkle proof checks (graph and weights) shrink monotonically with $N$ (roughly $2{\times}$$4{\times}$ reduction between N=2N{=}2 and $N{\geq}8$). Per-round substep latency: the rightmost panel shows the time for the proposer’s partition and the challenger’s selection at dispute round index $i$ within an $N=4$-way partition; both decay with $i$. Boxplot information is collected through eight dispute games across eight different perturbed operators through the model. Guideline and insight: $N\!\in\![8,12]$ minimizes wall-clock and on-chain work without creating oversized proofs.
7.3. End-to-end cost and on-chain footprint
Table 3. Forward and dispute costs across models ($N=2$). DCR is the challenger FLOP range to reach and adjudicate the leaf; Cost Ratio is DCR/forward FLOPs. Gas sums all dispute message interactions.
Metric BERT Diffusion LLM ResNet
Forward Cost (FLOPs, $10^{9}$) 19.47 802.49 485.09 23.09
Dispute Steps 11 12 13 11
On-chain Cost (k gas) 1984.4 2073.1 2161.8 1984.4
DCR (FLOPs, $10^{9}$) [9.74, 20.6] [316, 995] [223, 485] [11.7, 24.3]
Cost Ratio Range [0.50, 1.06] [0.39, 1.24] [0.46, 1.00] [0.51, 1.05]
[…]
Deterministic execution overhead. The deterministic settings introduced in Sec. 7.1 is also set during the optimistic execution phase (Phase 1 in Sec. 3.2) to ensure the system functions correctly. We test the overhead of these settings by running Qwen-8B model on 100 WikiText-103 (Merity et al., 2017) input data. The experiments show that the software deterministic settings introduce a negligible overhead of  0.3% in terms of latency compared to the non-deterministic settings.
[…]
Where NAO fits. When users accept tolerance-aware correctness, NAO offers orders-of-magnitude lower end-to-end latency and memory usage than zk proving for LLM-scale inference while preserving native kernels and heterogeneity. zk remains compelling when public verifiability with model secrecy is paramount; NAO is compelling when floating-point semantics and production performance are constraints.
8. Related Work
8.1. Verifiable Machine Learning
Work on verifiable ML splits into cryptographic validity proofs and optimistic approaches. Zero-knowledge systems compile models to finite-field circuits and attach succinct proofs (e.g., SafetyNets (Ghodsi et al., 2017), zkML (Chen et al., 2024), zkGPT (Qu et al., 2025), zkllm (Sun et al., 2024)), but still incurring substantial proving time/memory and often requiring quantization or restricted operators; TEE-assisted designs reduce proof cost but reintroduce trust in hardware vendors and exists performance degradation due to limited storage and compute capacity (Tramer and Boneh, [n. d.]; Moon et al., 2025; Bai et al., 2025; Niu et al., 2022). Optimistic protocols (Agatha (Zheng et al., 2021), opML (Conway et al., 2024)) and refereed delegation (Verde (Arun et al., 2025)) localize disputes on computation graphs and arbitrate on chain, with practicality often hinging on replication or enforcing bitwise-reproducible operators (RepOps) across heterogeneous accelerators; recent systems generalize dispute games to Turing-complete programs and explore incentive mechanisms such as Proof-of-Sampling (Mirkin et al., 2025; Zhang et al., 2024). NAO takes a complementary tack: rather than enforcing bitwise equality, it verifies results up to principled, operator-level tolerances that acknowledge nondeterminism. Unlike approaches that determinize execution (e.g., high-precision rounding logs for verifiable training or RepOps libraries (Srivastava et al., 2024; Arun et al., 2025)), NAO aims to preserve native GPU kernels and heterogeneity while still providing portable accountability.
8.2. Floating-Point Rounding Errors in Neural Networks
A growing body of work shows that ignoring floating-point effects renders verification unsound in practice: real-number proofs can certify networks that fail under IEEE-754 execution, and deployed stacks exhibit nondeterminism from non-associative reductions, kernel fusion, and scheduler variation (Jia and Rinard, 2021; Szász et al., 2025). Measurement tools such as FPRev reverse-engineer accumulation orders across software/hardware stacks(Xie et al., 2025). In response, some systems try to eliminate nondeterminism, e.g., reproducible operator libraries or training with high-precision states and logged rounding decisions to enable exact replay across GPUs, but at non-trivial performance and engineering cost (Arun et al., 2025; Srivastava et al., 2024).
9. Discussion and Conclusion
We introduce NAO, a tolerance-aware optimistic protocol that makes neural network computation verifiable on heterogeneous accelerators without relying on bitwise determinism or trusted hardware. NAO verifies results up to operator-specific acceptance regions by combining portable IEEE-754 bounds with tight empirical error percentile thresholds, and it localizes disputes to a single operator via a Merkle-anchored dispute game followed by a low-cost leaf check. This design preserves native kernels and performance while providing vendor-agnostic accountability; our implementation and attack study demonstrate the robustness of the tolerance-aware approach.
NAO naturally handles multi-step workloads (autoregressive decoding, diffusion sampling, and training) by layering time over the dispute game. Concretely, we commit to a temporal Merkle tree of step states (e.g., token logits, parameter checkpoints), bisect first across time to the earliest offending step, then within that step across the operator DAG to the leaf adjudication. This 2D partition yields prefix finality: early tokens/steps can finalize even while later steps remain subject to challenge. For training, we anchor checkpoints and optimizer state, verify per-update deltas, and permit windowed challenges over mini-batches. A minor limitation is calibration churn: onboarding new devices or kernels requires efficiently updating empirical thresholds and validating them. Extending NAO so that multi-step settings work by generalizing temporal bisection/$N$ way partition with windowed mini-batch challenges and streamlining threshold calibration is ongoing work.
Appendix A Floating Point
A.1. Floating Point Basics
We assume IEEE 754 arithmetic with rounding to nearest, ties to even, and exclude overflow/underflow/NaNs unless stated otherwise. Let $u$ denote the unit roundoff (machine epsilon divided by 2), e.g. $u=2^{-24}$ for float32. For any basic operation ${+,,×,/}\circ\in\{$$\times$,/\} on normalized operands, the standard model captures one rounding step as
(16) $\operatorname{fl}(x\circ y)\;=\;(x\circ y)(1+\delta),\qquad|\delta|\leq u.$
It is often convenient to write products of many $(1+\delta)$ terms as
(17) $\prod_{j=1}^{m}(1+\delta_{j})\;=\;1+\theta_{m},\qquad|\theta_{m}|\leq\gamma_{m}\quad\text{with}\quad\gamma_{m}\;\triangleq\;\frac{mu}{1-mu},$
valid for $mu<1$; see (Higham, 2002; Goldberg, 1991, Sec. 2–3). We will also use that $\gamma_{k}$ is monotone in $k$ and that $\gamma_{k}=ku+O(u^{2})$ for fixed $k$ and small $u$.
Modeling conventions. Unless stated otherwise, our bounds assume: (i) rounding to nearest, ties to even; (ii) no subnormal underflow/overflow; (iii) each rounding error $\delta$ obeys $|\delta|\leq u$ as in (16). When we invoke a probabilistic model, we additionally assume independent, mean-zero, bounded rounding errors as stated below.
A.2. Theoretical error bounds
In this section, we elaborate on the derivation of theoretical error bound formula for floating-point operations, including the deterministic and probabilistic bounds.
Let $u$ be the unit roundoff (e.g., $u=2^{-24}$ for float32). In the standard model (Higham, 2002; Goldberg, 1991),
$\operatorname{fl}(x\circ y)\;=\;(x\circ y)(1+\delta),\qquad|\delta|\leq u,\quad\circ\in\{+,-,\times,/\},$
For an $n$-term sum $s=\sum_{i=1}^{n}x_{i}$ evaluated by sequential FP adds,
$\widehat{s}_{k}=\operatorname{fl}(\widehat{s}_{k-1}+x_{k})=(\widehat{s}_{k-1}+x_{k})(1+\delta_{k}),\quad k=1,\dots,n,\quad\widehat{s}_{0}=0,$
one can unroll to obtain
(18) $\widehat{s}=\sum_{i=1}^{n}x_{i}\prod_{j=i+1}^{n}(1+\delta_{j}).$
Taking absolute values and using the triangle inequality,
(19) $|\widehat{s}-s|\;\leq\;\sum_{i=1}^{n}|x_{i}|\;\Big|\prod_{j=i+1}^{n}(1+\delta_{j})-1\Big|.$
We use the tail-product bound (Higham, 2002, Sec. 3.1): if $|\delta_{j}|\leq u$ and $mu<1$, then
(20) $\Big|\prod_{j=1}^{m}(1+\delta_{j})-1\Big|\;\leq\;\gamma_{m}\;\triangleq\;\frac{mu}{1-mu}.$
Applying the lemma to each tail with m=nim=n-i and using that $\gamma_{m}$ increases with $m$ yields the deterministic bound
(21) $|\widehat{s}-s|\;\leq\;\gamma_{n-1}\,\sum_{i=1}^{n}|x_{i}|,\qquad\text{valid when }(n-1)u<1.$
Following probabilistic model (Madry et al., 2018), assume the rounding errors in a length-$n$ computation are independent, mean-zero, and bounded by $u$. Then for any $\lambda>0$,
(22) $\prod_{i=1}^{n}(1+\delta_{i})\;$ = $=\;1+\theta_{n},$
(23) $|\theta_{n}|\;$ $\leq\;\widetilde{\gamma}_{n}(\lambda)\;\triangleq\;\exp\!\Big(\lambda\sqrt{n}\,u\;+\;\tfrac{nu^{2}}{1-u}\Big)-1,$
with probability at least
(24) $P(\lambda)\;=\;1-2\exp\!\Big(-\tfrac{\lambda^{2}(1-u)^{2}}{2}\Big).$
Since $e^{t}\leq\tfrac{1}{1-t}$ for $0\leq t<1$, we have
(25) $\widetilde{\gamma}_{n}(\lambda)\;\leq\;\lambda\sqrt{n}\,u+O(u^{2}),\qquad\lambda\sqrt{n}\,u+\tfrac{nu^{2}}{1-u}<1.$
Applied to summation (21), we obtain the high-probability bound
(26) $\Pr\!\left(|\widehat{s}-s|\;\leq\;\widetilde{\gamma}_{n-1}(\lambda)\sum_{i=1}^{n}|x_{i}|\right)\;\geq\;P(\lambda).$
A.3. Implemented PyTorch operation list for theoretical error bounds
We elaborate all the PyTorch operations for which we derive theoretical error bounds in both deterministic and probabilistic models. Operators include: basic arithmetic and elementwise functions (Add, Sub, Mul, Div, Pow, Neg, Sqrt, Rsqrt, Exp, Log, Sin, Cos, Tanh); activations (ReLU, GELU, SiLU); normalization and softmax (Softmax, BatchNorm, LayerNorm, GroupNorm); linear algebra and convolutions (Matmul, Bmm, Linear, Conv2D); reductions, pooling, upsampling (Mean, AdaptiveAvgPool2d, MaxPool2d, Interpolate); structural and non-arithmetic ops (Concatenate, Slice, Flatten, Indexing, Views/transforms, Type/device/movement, Masked fill, Embedding, Max/Min, Dropout eval).
Appendix B Stability of Empirical Error Percentile Profiles
Goal. We assess the stability of per-operator empirical error percentile profiles as additional samples accrue. Let $\mathcal{I}$ index operators, $\mathcal{P}$ the set of error percentiles (e.g., $p\in\{10,15,\dots,90\}$), and $n$ the sample size (e.g., $n=50$). For operator $i\in\mathcal{I}$ and percentile $p\in\mathcal{P}$, denote the per-sample sequence by $\{y_{i,p,t}\}_{t=1}^{n}$. The running median is
(27) $\widetilde{\theta}_{i,p}(k)\;=\;\operatorname{median}\big\{y_{i,p,1},\dots,y_{i,p,k}\big\},\qquad k=1,\dots,n,$
with $\theta_{i,p}:=\widetilde{\theta}_{i,p}(n)$ the point estimate at $n$. Unless stated otherwise, the median of an even number of points is the midpoint of the two central order statistics.
Relative scale. To obtain scale-free magnitudes we use the symmetric relative change
(28) $\delta(a,b)\;=\;\frac{2\,|a-b|}{|a|+|b|+\varepsilon},\qquad\varepsilon>0\ \text{small}.$
When normalizing a quantity $x$ by a baseline $z$ we use $|x|/(|z|+\varepsilon)$ with the same safeguard $\varepsilon$.
B.1. Diagnostics (per operator and per percentile)
We compute four complementary robustness diagnostics for each $(i,p)$, reported in relative units. Throughout, assume $W\in\{1,\dots,n-1\}$ and exclude non-finite values before computing any statistics.
  1. (D1)
    Short-horizon relative drift (SupNorm). Detects drift of the running median within the last $W$ steps:
    (29) $\mathrm{SupNorm}_{i,p}\;=\;\max_{k\in\{n-W,\dots,n-1\}}\delta\!\big(\widetilde{\theta}_{i,p}(n),\,\widetilde{\theta}_{i,p}(k)\big).$
    (This matches the usual supremum interpretation over the most recent horizon. If a single-lag version is preferred, replace the max by k=nWk=n-W.)
  2. (D2)
    Maximum leave-one-out influence (Jackknife). Measures sensitivity to any single sample:
    (30) $\mathrm{JK}_{i,p}\;=\;\max_{1\leq t\leq n}\;\frac{\big|\operatorname{median}(\{y_{i,p,s}\}_{s\neq t})-\theta_{i,p}\big|}{|\theta_{i,p}|+\varepsilon}.$
    (Optionally replace the numerator with $\delta\!\big(\operatorname{median}(\{y_{i,p,s}\}_{s\neq t}),\allowbreak\,\theta_{i,p}\big)$ for full symmetry.)
  3. (D3)
    Tail adjustment over the last $W$ steps (TailAdj). Detects lingering adjustments in the tail of the running-median curve:
    (31) $\mathrm{TailAdj}_{i,p}\;=\;\max_{k\in\{n-W,\dots,n-1\}}\frac{\big|\widetilde{\theta}_{i,p}(k\!+\!1)-\widetilde{\theta}_{i,p}(k)\big|}{|\theta_{i,p}|+\varepsilon}.$
  4. (D4)
    Rolling-window variability (RollSD). Captures local variability of windowed medians. Let $r_{i,p}(k)$ be the median of a length-$W$ sliding window ending at $k$ (defined for $k=W,\dots,n$). Report
    (32) $\mathrm{RollSD}_{i,p}\;=\;\frac{\operatorname{sd}\big(\{r_{i,p}(k)\}_{k=W}^{n}\big)}{|\theta_{i,p}|+\varepsilon}.$
B.2. Global drift across percentiles (per operator)
To summarize cross-percentile instability, each operator is represented by the worst-case short-horizon drift over the last $W$ steps:
(33) $\Delta_{i}^{(\infty)}\;=\;\max_{p\in\mathcal{P}}\ \max_{k\in\{n-W,\dots,n-1\}}\delta\!\big(\widetilde{\theta}_{i,p}(n),\,\widetilde{\theta}_{i,p}(k)\big).$
B.3. Aggregation and reporting
All diagnostics are computed per operator and per percentile and then aggregated across operators. For each percentile $p$ (i.e., each column abs_qpp), we report the median (50th) and upper-decile (90th) across operators.
Default settings. Unless stated otherwise, we use $W=10$ (tail/window length), $n=50$ samples per operator, and $\mathcal{P}=\{10,15,\dots,90\}$. Non-finite values are excluded prior to any per-operator summaries. All normalizations use (28) with the same $\varepsilon$.